OpenStudy (anonymous):

√(4-x)+ √(3x+100) = 0

OpenStudy (anonymous):

help me fin its solution set :)

OpenStudy (anonymous):

why not get √(4-x) on right hand side and then square both sides?

OpenStudy (anonymous):

√(3x+100) = -√(4-x) 3x+100= (-1*√(4-x))^2 3x+100=(-1)^2 * (√(4-x))^2 3x+100= 4-x i think.

OpenStudy (anonymous):

then I nedd to find the value of x

OpenStudy (anonymous):

yes but now it is more simple :)

OpenStudy (anonymous):

ok, thank you for helping me :)

OpenStudy (anonymous):

I could be wrong tho :(

OpenStudy (anonymous):

its ok :) I appreciate your help :)

OpenStudy (anonymous):

√(4-x)+ √(3x+100) = 4 correction :(

OpenStudy (anonymous):

its supposed to be four, not zero

OpenStudy (anonymous):

oh that makes it harder lol. Imma try

OpenStudy (anonymous):

LOL im not good at math T.T

OpenStudy (mathstudent55):

Is this the equation? \( \sqrt{4-x} + \sqrt{3x+100} = 4\)

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

fidn the solution set :)

OpenStudy (mathstudent55):

You will need to square it twice. First, you isolate one of the square roots, and you square both sides. Then you isolate the remaining square root, and you square both sides again. Finally, when you get answers for x, you check all of them in the original equation because squaring an equation introduces extraneous solution.

OpenStudy (anonymous):

okay...

OpenStudy (mathstudent55):

Here we go: \( \sqrt{4-x} + \sqrt{3x+100} = 4\) Isolate one root: \( \sqrt{4-x} = 4 - \sqrt{3x+100}\) Now when we square both sides, the left root will disappear, but we will still have the one on the right side to deal with.

OpenStudy (mathstudent55):

\( (\sqrt{4-x})^2 = (4 - \sqrt{3x+100})^2 \) \( 4 - x = 16 - 8\sqrt{3x+100} + 3x + 100 \) Do you follow so far?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

where did you get the -8?

OpenStudy (mathstudent55):

The right side was: \((4 - \sqrt{3x+100})^2 \) This is the square of a binomial which follows this pattern: \( (a - b)^2 = a^2 - 2ab + b^2 \) so \((4 - \sqrt{3x+100})^2 \) \( = 4^2 - 2(4)(\sqrt{3x+100}) + (\sqrt{3x+100})^2 \) There you see there the -8 comes from.

OpenStudy (anonymous):

ok, I remember that :)

OpenStudy (anonymous):

what do we do now?

OpenStudy (mathstudent55):

Let's continue from where we left off. We got to here: \(4 - x = 16 - 8\sqrt{3x+100} + 3x + 100 \)

OpenStudy (mathstudent55):

Now we need to again isolate the root, and move everything else to the other side. Then we square both sides.

OpenStudy (anonymous):

yessss :)

OpenStudy (mathstudent55):

I added the root to both sides, moving the root to the left side. Then I subtracted 4 from both sides moving the 4 to the right side. I added x to both sides moving the x to the right side. \( 8\sqrt{3x+100} = 16 + 3x + 100 -4 + x\) Now we combine like terms on the right side: \( 8\sqrt{3x+100} = 4x + 112\)

OpenStudy (mathstudent55):

Now we have the root n the left side by itself, so we square both sides again: \( (8\sqrt{3x+100})^2 = (4x + 112)^2 \) \(64(3x + 100) = (4x)^2 + 2(4x)112 + 112^2 \)

OpenStudy (anonymous):

okay :)

OpenStudy (mathstudent55):

\(192x + 6400 = 16x^2 + 896x + 12,544 \)

OpenStudy (mathstudent55):

\(0 = 16x^2 + 704x + 6144 \)

OpenStudy (anonymous):

factor it :)

OpenStudy (mathstudent55):

Exactly. Now we divide both sides by 16 and switch sides: \(x^2 + 44x + 384 = 0 \)

OpenStudy (anonymous):

whooops sorry

OpenStudy (mathstudent55):

Now we need to either factor the quadratic or use the quadratic formula.

OpenStudy (anonymous):

its easier if we will use the quadratic formula.

OpenStudy (mathstudent55):

Exactly what I was thinking. Let's use the quadratic formula. \( x = \dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a } \) Here, we have \(a = 1\), \(b = 44\), \(c = 384\).

OpenStudy (mathstudent55):

\( x = \dfrac{-44 \pm \sqrt{44^2 - 4(1)(384)}}{2(1) } \) \( x = \dfrac{-44 \pm \sqrt{1936 - 1536}}{2 } \) \( x = \dfrac{-44 \pm \sqrt{400}}{2 } \) \( x = \dfrac{-44 \pm 20}{2 } \) \(x = -12\) or \(x = -32\)

OpenStudy (mathstudent55):

Ok, so far?

OpenStudy (anonymous):

ok :) so the solution set is the set containing -12 and -32

OpenStudy (mathstudent55):

Not so fast.

OpenStudy (anonymous):

yehey! thank you so much for helping me :)

OpenStudy (anonymous):

??????

OpenStudy (mathstudent55):

Way above, when I told you the steps you need to follow to solve this equation I wrote: "You will need to square it twice. First, you isolate one of the square roots, and you square both sides. Then you isolate the remaining square root, and you square both sides again. Finally, when you get answers for x, you check all of them in the original equation because squaring an equation introduces extraneous solution."

OpenStudy (mathstudent55):

Read the last instruction, starting with "Finally, ..."

OpenStudy (anonymous):

ok, so its an extraneous solution

OpenStudy (mathstudent55):

I didn't say that. I just said we need to test both soltions we got in the original equation to see if they are solutions or not.

OpenStudy (anonymous):

ah okay :)

OpenStudy (anonymous):

its the solution

OpenStudy (anonymous):

4+8=12 :)

OpenStudy (mathstudent55):

1. Let's test x = -12 \(\sqrt{4-(-12)} + \sqrt{3(-12)+100} =?~ 4\) \(\sqrt{16)} + \sqrt{64} =?~ 4\) \(4 + 8 =? ~ 4 \) \(12 \ne 4 \) x = -12 does not work.

OpenStudy (anonymous):

? im wrong :(

OpenStudy (mathstudent55):

1. Let's test x = -32 \(\sqrt{4-(-32)} + \sqrt{3(-32)+100} =?~ 4\) \(\sqrt{36)} + \sqrt{4} =?~ 4\) \(6 + 2 =? ~ 4 \) \(8 \ne 4 \) x = -32 does not work.

OpenStudy (anonymous):

OMG

OpenStudy (mathstudent55):

Neither solution worked, so the answer is there is no solution. As a set, the answer is { }, the empty set.

OpenStudy (anonymous):

omg, its empty set LOL sorry

OpenStudy (anonymous):

I failed in my math exam because I was careless :(

OpenStudy (mathstudent55):

Don't be sorry. This was a good example of how squaring both sides of an equation can introduce extraneous solutions. The solutions we got were correct solutions to the quadratic equation, but the quadratic equation did not exactly correspond to the original equation. That's why solutions to the quadratic equation were not solutions to the original equation.

OpenStudy (mathstudent55):

I made plenty of careless mistakes myself. You just need to check your work.

OpenStudy (anonymous):

extraneous...big word :) OK, I will remember that :)

OpenStudy (anonymous):

thank you :) hahaha

OpenStudy (anonymous):

im a neophyte here, how can I give you a medal? :)

OpenStudy (mathstudent55):

Make sure you remember what it means. In our case, the extraneous solutions were false solutions that were introduced by squaring both sides. That's why it's important to check any possible solutions in the original equation.

OpenStudy (anonymous):

okay :) my professor in math mentioned that

OpenStudy (mathstudent55):

Click the blue "Best Reposne" box in one of my responses.

OpenStudy (anonymous):

done? hahaha thanks :)

OpenStudy (mathstudent55):

wlcm